OpenStudy (anonymous):

ln(x + 4) − ln(x) = 2

OpenStudy (anonymous):

\[\ln(\frac{x+4}{x})=2\] \[\frac{x+4}{x}=e^2\] \[x+4=e^2x\] \[4=e^2x-x\] \[4=(e^2-1)x\] \[x=\frac{4}{e^2-1}\]